Solid ethyl alcohol, 30 g, at -180.3 oC is vaporized at the boiling point of 78.3 oC.
Cp (solid) = 0.971 J/(g*K), Cp (liquid) = 2.30 J/(g*K), Tfreezing = -117.3 oC, ∆fusH = 218 J/g, ∆vapH =
854 J/g
62-64) How much heat was involved from initial state up to end of melting?
65-67) How much heat was involved from right upon melting up to reaching boiling point?
68-70) How much heat was involved in the whole process?
1
Expert's answer
2020-11-20T08:09:42-0500

62-64).

Q= mc∆T.

Q1= 30 x 0.971 x 180.3 =5 252,139 J.

Q2= 30 x 2.3 x 78.3 = 5 402,7 J.

Q3= 30 x 218 x 78.3 = 512 082 J.

Qtotal= Q1+Q2+Q3 = 5252,139+5402,7+512082 = 522 736,839 J.

65-67). Q2= 30 x 2.3 x 78.3 = 5 402,7 J.

68-70).

522 736,839 J.